OpenStudy (anonymous):

find the intersection points of x^2+y^2=4 and y=x-2

OpenStudy (taylor<3srin):

First put the first equation into Slope Intercept form.

OpenStudy (anonymous):

y=x+2 y=x-2

OpenStudy (taylor<3srin):

Those two lines dont intersect do they?

OpenStudy (anonymous):

lol i meant y=-x+2 and y=x-2

OpenStudy (taylor<3srin):

Since I am tired, and don't feel like explaining the whole thing, the points intersect at (2,0)

OpenStudy (anonymous):

lol thanks taylor

OpenStudy (taylor<3srin):

No problemo seniorita